OpenStudy (anonymous):

Find the radius of a circle with the given circumference. 12pi in.

OpenStudy (anonymous):

Well, circumference =2pi(r) 12pi=2pi(r) 12pi/2=2pi(r)/2 6pi=pi(r) 6pi/pi=pi(r)/pi 6=(r) So, the radius is 6in.